Production Budget Assume that Stillwater Designs produces two automotive subwoofers: S12L7 and S12L5. The S12L7 sells for $475, and the S12L5 sells for $300. Projected sales (number of speakers) for the coming five quarters are as follows: S12L7 S12L5 First quarter, 20Y1 880 1,430 Second quarter, 20Y1 2,420 1,540 Third quarter, 20Y1 6,160 5,830 Fourth quarter, 20Y1 5,060 4,290 First quarter, 20Y2 990 1,320 The vice president of sales believes that the projected sales are realistic and can be achieved by the company. Stillwater Designs needs a production budget for each product (representing the amount that must be outsourced to manufacturers located in Asia). Beginning inventory of S12L7 for the first quarter of 20Y1 was 340 boxes. The company's policy is to have 20% of the next quarter's sales of S12L7 in ending inventory. Beginning inventory of S12L5 was 170 boxes. The company's policy is to have 30% of the next quarter's sales of S12L5 in ending inventory. Required: Prepare a production budget for each quarter for 20Y1 and for the year in total. Stillwater Designs Production Budget for S12L7 For the Year Ended December 31, 20Y1 1st Qtr. 2nd Qtr. 3rd Qtr. 4th Qtr. Year Sales Desired ending inventory Total needs Less: Beginning inventory Units produced Prepare a production budget for each quarter for 20Y1 and for the year in total. If required, round your answers to nearest whole value. Stillwater Designs Production Budget for S12L5 For the Year Ended December 31, 20Y1 1st Qtr. 2nd Qtr. 3rd Qtr. 4th Qtr. Year Sales Desired ending inventory Total needs Less: Beginning inventory Units produced
Answer #1

Answer:

Production budget for each quarter for 20Y1 and for the year in total for S12L7 is as follows:

Stillwater Designs Production Budget for S12L7 For the Year Ended December 31, 20Y1 2011 1st Qtr. 2nd Qtr 3rd Qtr. Sales 880

As given:

Ending inventory = 20% * Sales of next Quarter

Hence: Ending inventory of Q4 = 990 * 20% = 198

Beginning inventory of a quarter = Ending inventory of previous quarter

Beginning inventory of Q1, as given = 340

Production budget for each quarter for 20Y1 and for the year in total for S12L5 is as follows:

Stillwater Designs Production Budget for S712L5 For the Year Ended December 31, 2041 2011 Q1 Q2 Q3 Sales Units 1,430 1,540 5,

As given:

Ending inventory = 30% * Sales of next Quarter

Hence: Ending inventory of Q4 = 1320 * 30% = 396

Beginning inventory of a quarter = Ending inventory of previous quarter

Beginning inventory of Q1, as given = 170